# Enhancing Accuracy-Privacy Trade-off in Differentially Private Split Learning

### v2

@@ I Introduction
Unfortunately, merely sharing smashed data does not provide an absolute guarantee of protecting raw private data. The shared smashed data inherently contains substantial latent information about the input data and could potentially serve as a basis for relatively straightforward reversal techniques, such as visualization [6, 7], or more intricate reconstructions like model inversion attacks [8, 9, 10]. Several efforts [6, 7, 8, 11, 12, 13, 14] aim to mitigate this risk through the implementation of differential privacy (DP), which is currently a standard practice in privacy-preserving machine learning. For example, in [14], the researchers explore the utilization of DP to safeguard users’ data privacy during collaborative inference using SL for IoT. Fig. 1 visually illustrates the integration of DP noise into smashed data to enhance data privacy preservation in the context of SL. A notable privacy vulnerability of SL arises from the fact that a neural network operates as a differentiable smooth function, inherently susceptible to functional inversion. Partitioning the model between clients and the server does not alter this characteristic [15]. Researchers propose countering this vulnerability by introducing DP noise to the smashed data, making individual identification no longer feasible. They assess the effectiveness of their approach across various image datasets, demonstrating that DP can effectively safeguard the privacy of users data while achieving a degree of accuracy reduction.
@@
In much of the existing literature, it is noteworthy that calibrated noise is primarily added exclusively to the outputs from the split layer of SL, with little consideration to applying DP to other local layers. In [8], alongside their comprehensive privacy attack proposals, the authors explore the application of noise at the original input to defense, claiming its superior effectiveness compared to noise at the split layer. However, this claim lacks a deeper investigation or thorough evaluation. Nevertheless, a central challenge with DP lies in striking the right balance between preserving model utility and ensuring data privacy. Increased noise levels can provide stronger privacy guarantees, but inevitably impact the overall model utility (accuracy). Concerning SL with DP, a fundamental question arises: Within the SL framework, where should DP be applied to achieve the optimal trade-off between model accuracy and data privacy? To address this research question, we examine the trade-offs between accuracy and privacy when applying DP at different local layers of SL to identify the most effective position for introducing noise and achieving an improved trade-off. Our findings indicate that placing DP noise farther from the input layer yields a superior balance. Moreover, introducing noise to the input layer helps to reduce data leakage more effectively than noise in other layers but significantly reduces model accuracy, leading to an unfavorable trade-off. Based on these insights, we propose a method to reduce the exposed (smashed) data size, limiting data leakage and reducing SL communication costs while maintaining the most favorable accuracy and privacy trade-offs. In summary, our contributions in this paper are as follows:
@@ II-C SL with differential privacy
$\mathcal{F}$ is said to provide $(\epsilon,\delta)$-DP, with $s$ representing the sensitivity of $f$. Our choice of the Gaussian mechanism is motivated by its ability to accommodate the summation of two distributions, which will be utilized in our subsequent proposal. Considering $f$ as the local model, the literature encompasses a substantial body of work [6, 7, 8, 11, 12, 13, 14] dedicated to introducing DP noise to smashed data to mitigate the risk associated with raw data reconstruction. Furthermore, based on the post-processing property, DP noise can be dynamically applied to the split layer, input, or any local layer without violating the privacy guarantee, as shown in Eq. 4 below:
@@ III-A Different privacy requirements
The clients require DP by adding Gaussian noise to their smashed data before transmitting it to the server. Within each training round, the order of client training with the server is randomly shuffled. The post-training accuracy results are summarized in Table I, revealing that incorporating the DP noise of some clients leads to a decrease in training accuracy for all clients compared to scenarios where all clients demand the same DP noise. In particular, when no client requires DP, the accuracy is $99.2\%$, which is slightly reduced to $98.8\%$ in the scenario of only one client is requiring DP. Regarding the DP-enabled client, its accuracy is $66.3\%$, which is significantly impacted compared to the averaged value of $73.4\%$ when all clients require the same DP noise. The uniform privacy requirement, which allocates an equal privacy budget to each client, optimally facilitates achieving the highest accuracy regarding their respective noise levels. Thus, we consider the learning outcomes under uniform privacy as the benchmark upper limit for assessing accuracy loss or improvement in the context of varying privacy requirements. From the results, it is observed that DP-enabled clients are significantly affected while others are less so, which can lead to reluctance in contributing data from those clients with high privacy requirements. The experimental observations offer insights into what could be termed the ‘forgetting phenomenon’ [18] of SL, in which the server’s learning process on smashed data is sequentially impacted by different distributions (smashed data with and without additional DP noise). In situations where the majority of clients require no noise, the server might ‘forget’ the noise distribution while learning with ‘clean’ smashed data. This ultimately diminishes accuracy when working with noisy smashed data from DP-enabled clients. Therefore, in the subsequent section, a novel training methodology is proposed to address this challenge.
@@ IV Utility and Privacy Trade-offs in SL with DP
While neural network architectures vary widely across applications, they generally consist of conceptually simple units called neurons. Each neuron computes a weighted sum of its inputs along with a bias term, $\mathbf{z}=\mathbf{w}\cdot\mathbf{x}+b$, followed by an element-wise nonlinear transformation, $\mathbf{a}=\phi(\mathbf{z})$ [23]. Due to the post-processing property of DP, privacy is still guaranteed whether noise is introduced to the layer’s output, $\mathbf{a}+\mathcal{N}$, or to the input of the layer, $\phi\left(\mathbf{w}\cdot(\mathbf{x}+\mathcal{N})+b\right)$. However, noise applied to the input data of a basic neuron block can be amplified due to the weight (e.g., $\mathbf{w}\cdot\mathcal{N}$) or the non-linearity (e.g., $\phi(\mathbf{z}+\mathbf{w}\cdot\mathcal{N})$), potentially having a higher impact on model utility. Drawing inspiration from [8], which investigates the effects of adding noise to input or split layers to mitigate data leakage from reconstruction attacks, we empirically analyze the application of noise across different local layers of SL in this section. Currently, there is a lack of theoretical analysis on the accuracy-privacy trade-offs when introducing noise across local layers of SL. Our empirical investigation aims to highlight the potential for such theoretical analysis. We begin by defining the threat model for our data leakage analysis, and then present by our findings on the optimal position for introducing noise to achieve the most favorable trade-offs between privacy and accuracy.
@@ VI Conclusion
This paper investigates the delicate balance between accuracy and privacy when leveraging DP to protect data privacy in SL. The study primarily explores the trade-offs between accuracy and privacy by introducing DP noise to different SL local layers with varying privacy budgets. Furthermore, based on the insights into the accuracy and privacy trade-offs, a method is proposed to reduce the size of smashed data, aiming to limit the data leakage and lower communication overhead, which is one of the notable drawbacks of SL. Expanding the scope to encompass a multi-client environment, with different privacy requirements, reveals potential reductions in accuracy among clients following training. To tackle the identified challenge, the study proposes a novel approach to enhance the learning performance by allowing the server to review clients’ noise distributions during training. Through a comprehensive array of experiments utilizing real-world datasets, the results presented demonstrate a significant improvement in accuracy for all clients requiring noise, effectively aligning their performance with the benchmark reference. Examining the accuracy-privacy trade-off and introducing strategies to enhance client accuracy provide valuable insights for effectively preserving data privacy in the context of distributed learning across heterogeneous devices, such as mobiles and IoT. Future research directions include exploring the optimal augmented data size to optimize server computations, examining on other learning tasks, network models, and privacy-preserving techniques.

### v3

@@ Abstract
Index Terms:
Multi-client split learning, differential privacy, privacy preservation, data leakage, reconstruction attacks.
@@ I Introduction
Unfortunately, merely sharing smashed data does not provide an absolute guarantee of protecting raw private data. The shared smashed data inherently contains substantial latent information about the input data and could potentially serve as a basis for relatively straightforward reversal techniques, such as visualization [6, 7], or more intricate reconstructions like model inversion attacks [8, 9, 10]. Several efforts [6, 7, 8, 11, 12, 13, 14, 15, 16] aim to mitigate this risk through the implementation of differential privacy (DP), which is currently a standard practice in privacy-preserving machine learning. For example, in [14], the researchers explore the utilization of DP to safeguard users’ data privacy during collaborative inference using SL for IoT. Fig. 1 visually illustrates the integration of DP noise into smashed data to enhance data privacy preservation in the context of SL. A notable privacy vulnerability of SL arises from the fact that a neural network operates as a differentiable smooth function, inherently susceptible to functional inversion. Partitioning the model between clients and the server does not alter this characteristic [17]. Researchers propose countering this vulnerability by introducing DP noise to the smashed data, making individual identification no longer feasible. They assess the effectiveness of their approach across various image datasets, demonstrating that DP can effectively safeguard the privacy of users data while achieving a degree of accuracy reduction.
@@
In much of the existing literature, it is noteworthy that calibrated noise is primarily added exclusively to the outputs from the split layer of SL, with little consideration to applying DP to other local layers. In [8], alongside their comprehensive privacy attack proposals, the authors explore the application of noise at the original input to defense, claiming its superior effectiveness compared to noise at the split layer. However, this claim lacks a deeper investigation or thorough evaluation. Nevertheless, a central challenge with DP lies in striking the right balance between preserving model utility and ensuring data privacy. Increased noise levels can provide stronger privacy guarantees, but inevitably impact the overall model utility (accuracy). Concerning SL with DP, a fundamental question arises: Within the SL framework, where should DP be applied to achieve the optimal trade-off between model accuracy and data privacy? To address this research question, we examine the trade-offs between accuracy and privacy when applying DP at different local layers of SL to identify the most effective position for introducing noise and achieving an improved trade-off. Our findings suggest that applying DP noise further from the input layer achieves a better balance between privacy and accuracy. Conversely, at the same DP level, adding noise directly to the input layer more effectively reduces data leakage but significantly compromises model accuracy, resulting in a less favorable trade-off. Based on these insights, we propose a method to reduce the exposed (smashed) data size, limiting data leakage and reducing SL communication costs while maintaining the most favorable accuracy and privacy trade-offs. In summary, our contributions in this paper are as follows:
@@ II-C SL with differential privacy
$\mathcal{F}$ is said to provide $(\epsilon,\delta)$-DP, with $s$ representing the sensitivity of $f$. Our choice of the Gaussian mechanism is motivated by its ability to accommodate the summation of two distributions, which will be utilized in our subsequent proposal. Considering $f$ as the local model, the literature encompasses a substantial body of work [6, 7, 8, 11, 12, 13, 14, 15, 16] dedicated to introducing DP noise to smashed data to mitigate the risk associated with raw data reconstruction. Additionally, due to the post-processing property of DP [23], noise can be applied dynamically to the split layer, input, or any local layer without compromising the privacy guarantees, as shown in Eq. 4 below:
@@ III-A Different privacy requirements
The clients ensure $(\epsilon,\delta)$-DP by adding Gaussian noise to their smashed data, as outlined in Eq. 3, before transmitting it to the server. Within each training round, the order of client training with the server is randomly shuffled. The post-training accuracy results are summarized in Table I, revealing that incorporating the DP noise of some clients leads to a decrease in training accuracy for all clients compared to scenarios where all clients demand the same DP noise. In particular, when no client requires DP, the accuracy is $99.2\%$, which is slightly reduced to $98.8\%$ in the scenario of only one client is requiring DP. Regarding the DP-enabled client, its accuracy is $66.3\%$, which is significantly impacted compared to the averaged value of $73.4\%$ when all clients require the same DP noise. The uniform privacy requirement, which allocates an equal privacy budget to each client, optimally facilitates achieving the highest accuracy regarding their respective noise levels. Thus, we consider the learning outcomes under uniform privacy as the benchmark upper limit for assessing accuracy loss or improvement in the context of varying privacy requirements. From the results, it is observed that DP-enabled clients are significantly affected while others are less so, which can lead to reluctance in contributing data from those clients with high privacy requirements. The experimental observations offer insights into what could be termed the ‘forgetting phenomenon’ [20] of SL, in which the server’s learning process on smashed data is sequentially impacted by different distributions (smashed data with and without additional DP noise). In situations where the majority of clients require no noise, the server might ‘forget’ the noise distribution while learning with ‘clean’ smashed data. This ultimately diminishes accuracy when working with noisy smashed data from DP-enabled clients. Therefore, in the subsequent section, a novel training methodology is proposed to address this challenge.
@@ IV Utility and Privacy Trade-offs in SL with DP
While neural network architectures vary widely across applications, they generally consist of conceptually simple units called neurons. Each neuron computes a weighted sum of its inputs along with a bias term, $\mathbf{z}=\mathbf{w}\cdot\mathbf{x}+b$, followed by an element-wise nonlinear transformation, $\mathbf{a}=\phi(\mathbf{z})$ [26]. Due to the post-processing property of DP, privacy is still guaranteed whether noise is introduced to the layer’s output, $\mathbf{a}+\mathcal{N}$, or to the input of the layer, $\phi\left(\mathbf{w}\cdot(\mathbf{x}+\mathcal{N})+b\right)$. However, noise applied to the input data of a basic neuron block can be amplified due to the weight (e.g., $\mathbf{w}\cdot\mathcal{N}$) or the non-linearity (e.g., $\phi(\mathbf{z}+\mathbf{w}\cdot\mathcal{N})$), potentially having a higher impact on model utility. Drawing inspiration from [8], which investigates the effects of adding noise to input or split layers to mitigate data leakage from reconstruction attacks, we empirically analyze the application of noise across different local layers of SL in this section. Currently, there is a lack of theoretical analysis on the accuracy-privacy trade-offs when introducing noise across local layers in SL. Our empirical investigation seeks to underscore the potential for such analysis. We begin by defining the threat model for our data leakage analysis, and then present by our findings on the optimal position for introducing noise to achieve the most favorable trade-offs between privacy and accuracy.
@@ VI Conclusion
This paper investigates the delicate balance between accuracy and privacy when leveraging DP to protect data privacy in SL. The study primarily explores the trade-offs between accuracy and privacy by introducing DP noise to different SL local layers with varying privacy budgets. Furthermore, based on the insights into the accuracy and privacy trade-offs, a method is proposed to reduce the size of smashed data, aiming to limit the data leakage and lower communication overhead, which is one of the notable drawbacks of SL. Expanding the scope to encompass a multi-client environment, with different privacy requirements, reveals potential reductions in accuracy among clients following training. To tackle the identified challenge, the study proposes a novel approach to enhance the learning performance by allowing the server to review clients’ noise distributions during training. Through a comprehensive array of experiments utilizing real-world datasets, the results presented demonstrate a significant improvement in accuracy for all clients requiring noise, effectively aligning their performance with the benchmark reference. Examining the accuracy-privacy trade-off and introducing strategies to enhance client accuracy provide valuable insights for effectively preserving data privacy in the context of distributed learning across heterogeneous devices, such as mobiles and IoT. Future research directions include conducting formal analysis, exploring the optimal augmented data size to optimize server computations, examining other learning tasks, network models, and privacy-preserving techniques.